< Chapter 1 Quiz Exercise 1.120 Part A Use the denity of aluminum to find the radus of the sphere in inches in 0
Answer #1

Density of Al = 2.70 g/cm3

Means,

Volume of 2.70 g. of Al = 1 cm3

then,

Volume of 85 g. of Al = 85 * 1/ 2.70 = 31.5 cm3

But,

1 inch = 2.54 cm

1 cm = 1 / 2.54 inch

1 cm3 = ( 1 / 2.54 ) 3 inch3

1 cm3 = 0.0610 inch3

Then, 31.5 cm3 = 0.0610 * 31.5 = 1.92 inch3.